9. Let y= X2, where X has the pdf below (a) Find the mean of Y without finding the pdf of Y. (b) Write the pdf of Y:f<v). (c) Find the mean of Y using.ffy), confirming your answer in part (a).
Answer #1

a)

mean of Y =E(Y)=E(X2)=\int_{0}^{\infty }x2 f(x) dx=\int_{0}^{\infty }x3e-x dx=-x3e-x-3x2e-x-6xe-x-6e-x |\infty0 =6

b)

here as pdf of Y =|dx/dy|fx(g-1(y))

for X=√Y

therefore |dx/dy| =1/(2√y))

hence pdf of Y f(y)= (1/(2√y)))*(√y*e-√y) =(1/2)*e-√y for 0 <y<\infty

c)

mean of Y =E(Y)=\int_{0}^{\infty }y*f(y) dy =\int_{0}^{\infty }(1/2)ye-√y dy

letting √y=u

dy=2√y du =2 u du

hence E(Y)=\int_{0}^{\infty }u3e-u du   =-u3e-u-3u2e-u-6ue-u-6e-u |\infty0 =6